A device contains three components, each of which is equally likely to be working or broken at any given time.

a) What is the probability that all three components are broken?

b) If at least two components need to be working for the device to function, what is the probability that the device functions?

c) Given that the device functions, what is the probability that all three components are working?
Answer #1

a)probability that all three components are broken =0.53 =0.125

b) probability that the device functions =P(2 or 3 component working =3C2(0.5)2(0.5)+3C3(0.5)3(0.5)0 =0.5

c)probability that all three components are working given device functions

=P(three component working)/P(at least two are working)=0.125/0.5 =0.25